1. Explain the principle involved in the separation of components of a mixture of liquids by using fractionating
column. What industrial applications does this process have?
2.The specific rotation of pure (R)-2-butanol is -13.5°. What % of a mixture of the two enantiomeric forms is (S)-2-butanol if the specific rotation of this mixture is -5.4° ?
3.Bromination of alkanes is a much slower reaction than chlorination.what of
expected to be the major organic product when 2-methylbutane is allowed to react with Br2 in the presence of light or heat?
1
Expert's answer
2020-06-08T15:04:31-0400

1) The components of the mixtures usually have different boiling points. Due to this, they can be separated by converting the mixture into a gaseous state and collecting the components according to their boiling points. Example of application in industry - primary oil refining. The primary end-products produced in petroleum refining may be grouped into four categories: light distillates(Liquified petroleum gas (LPG) and gasoline), middle distillates(kerosene and diesel fuels), heavy distillates (heavy fuel oils, wax, lubricating oils, asphalt) and others.


2) Excess of (S)-2-butanol =100% * 5.4 / 13.5 = 40%, so % (S)-2-butanol = (% excess of (S)-2-butanol + 1)/2 = (0.4+1)/2 = 0.7. We have a mixture of 70% (S)-2-butanol and 30% (R)-2-butanol.


3) Bromination of alkanes by heating or lighting takes place by a radical mechanism: